Lucy has invested some money invested at 18% and three times as much invested at 12%. her total annual income from the 2 investment is 85,000 pesos. how much was invested at each rate?
Answer by checkley77(12844)

.18x+.12*3x=85,000
.18x+.36x=85,000
.54x=85,000
x=85,000/.54
x=157,407.40 invested @ 18%
3*157,407.40=472,222.20 invested @ 12%
proof:
.18*157,407.40+.12*472,222.20=85,000
28,333.33+56,666.67=85,000
85,000=85,000